OpenStudy (anonymous):

find set of parametric equations for the line which passes through the point (3,4,7) that is perpendicular to the line with parametric equations x=1+t, y=3-t, and z=2+3t

OpenStudy (anonymous):

x=1+t, y=3-t, and z=2+3t {1,-1,3}.{a,b,c}=0 a-b+3c=0 {2,-1,-1} x=3+2t y=4-t z=7-t

OpenStudy (anonymous):

oh, set of ... a=b-3c {b-3c,b,c} x=3+(b-3c)t y=4+bt z=7+ct

OpenStudy (anonymous):

thank you so much